4. Rolling Two Dice: If two dice are rolled one time, find the probability of 

getting these results. 

a. A sum of 9 

b. A sum of 7 or 11 

c. A sum greater than or equal to 10


1
Expert's answer
2022-06-07T12:44:48-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 & 1 & 2 & 3 & 4 & 5 & 6 \\ \hline
 1 & 2 & 3 & 4 & 5 & 6 & 7 \\
 \hdashline
 2 & 3 & 4 & 5 & 6 & 7 & 8 \\
 \hdashline
 3 & 4 & 5 & 6 & 7 & 8 & 9 \\
 \hdashline
 4 & 5 & 6 & 7 & 8 & 9 & 10 \\
 \hdashline
 5 & 6 & 7 & 8 & 9 & 10 & 11 \\
 \hdashline
 6 & 7 & 8 & 9 & 10 & 11 & 12 \\
 \hdashline
\end{array}"

a.


"P(Sum=9)=\dfrac{4}{36}=\dfrac{1}{9}"

b.


"P(Sum=7\ or\ Sum=11)=P(Sum=7)"

"+P(Sum=11)=\dfrac{6}{36}+\dfrac{2}{36}=\dfrac{2}{9}"

c.


"P(Sum\ge10)=P(Sum=10)+P(X=11)"

"+P(X=12)=\dfrac{3}{36}+\dfrac{2}{36}+\dfrac{1}{36}=\dfrac{1}{6}"